OpenStudy (anonymous):

Check Work ? Let f(x)=(x-6)(6x+4) Find the equation for the tangent line to the graph of f at x=1 Tangent line: y= ? My work..... f'(x)=(x-6)(6)+(1)(6x+4) I applied product rule =12x-32 f'(1)=12(1)-32 = -20 y-(-50)=-20(x-1) [the -50 I got from plugging 1 into orig eq.] y=-20x-30 ?

OpenStudy (whpalmer4):

Let's check by multiplying the expression for \(f(x)\) and taking the derivative of that. \[f(x) = (x-6)(6x+4) = 6x^2+4x-36x-24 = 6x^2-32x-24\]\[\frac{df(x)}{dx} =2*6x^{2-1}-1*32x^{1-1} = 12x-32\]So far so good. \[f'(1) = 12(1)-32 = -20\] \[f(1) = (1-6)(6(1)+4) = -5*10 = -50\]So we have to run our tangent line through the point \((1,-50)\) Using point-slope formula:\[y-y_0 = m(x-x_0)\]\[y-(-50) = -20(x-1)\]\[y+50 = -20x+20\]\[y=-20x-30\] I agree with your answer. Is it possible that you are supposed to provide it in different form? 100% certain you copied the problem correctly?

OpenStudy (anonymous):

This makes sense I must've misread the question or put the wrong answer. Thank you for helping me out :) and for the detailed explanation !

OpenStudy (whpalmer4):

You're welcome!